### v1

@@ 2.3 A unified picture
For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm a}/\tau_{\rm e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime). As the relative damping weakens (Eq. 22), the fixed point becomes unstable, and the system’s trajectory begins to oscillate around the fixed point. It ultimately settles into a stable limit cycle (overstable trap). The limit cycle expands as $\tau_{\rm a}/\tau_{\rm e}$ continues to decrease. When the trajectory crosses the separatrix (Eq. 23), the system becomes dynamically unstable. In this regime, the oscillations grow over time and planets eventually leave resonance (escape). In contrast, for the case of a small outer planet shown in Figure 6, all traps are stable.
Figure 7: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $0.1$. Blue points indicate cases where the planets are not trapped by the resonance, green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate that the planets eventually escape from the resonance. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{a}\sim 10^{5}\ \mathrm{yr}$ and $\tau_{a}/\tau_{e}\sim 3\times 10^{2}$. While the analytical model predicts overstable trapping in this regime, the simulations show escape from resonance. We hypothesize that as migration becomes faster, the perturbations grow stronger and become nonlinear, leading to amplified oscillation amplitudes and an increased likelihood of escaping from resonance.
Figure 8: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $10$. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$. While the analytical model predicts resonance trapping in this regime, the simulations show resonance crossing. We hypothesize that the discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect.
Figure 9: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q=1$. We compare our criteria with previous studies.We cite the trapping criterion from Huang & Ormel (2023) and the stable and escape criteria from Goldreich & Schlichting (2014). The solid lines represent our criteria, the dash lines represent the criteria cited from previous studies. In the case of equal-mass planets, it is evident that our criteria align more closely with the simulation results.
3 Simulations
In this section, we conduct numerical simulations for $2:1$ resonance to verify the different regimes and criteria analytically derived in the previous section. The numerical setup is described in Sec. 3.1. The comparison between simulations and analysis at two extreme mass ratios is presented in Sec. 3.2, and the extrapolation for any mass ratio is discussed in Sec. 3.3.
3.1 Numerical setup
We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012). The simulations are confined in conplanar circumstances, and the accelerations for orbital decay and eccentricity damping are given by
${a_{\rm m}}=-\frac{\textbf{v}}{2\tau_{a}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}\cdot\textbf{r})\cdot\textbf{r}}{r^{2}\tau_{e}},$ (27)
which are implemented through REBOUNDx (Tamayo et al., 2020).
Initially, we place the inner and outer planets at $1$ and $1.7$ au, respectively, making them slightly away from the $2$:$1$ MMR. Both planets have circular orbits with random phase angles. To mimic convergent migration, we only let the outer planet undergo inward migration on a time-scale of $\tau_{\rm a}$, whereas the eccentricities of both planets are damped on a time-scale of $\tau_{\rm e}$. Simulations are terminated at $t{=}2\tau_{\rm a}$, and the final outcomes of the planets in each run are determined as described in Sec. 2.3.
We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$). We vary different sets of $\tau_{\rm a}$ and $\tau_{\rm e}$ in each simulations.
3.2 Comparison at two extreme mass ratios
The dynamical outcomes of planet pairs that undergo convergent migration are shown in Figure 6. The results are dependent on the mass ratio of the planets and disk properties ($\tau_{\rm a}$ and $\tau_{\rm e}$). Basically four regimes are presented when the inner planet is lower than that of the outer one. The illustrated runs in each regime are provided in Figures 1,2,3 and 5.
Figure 8 shows the numerical results obtained from an illustrated case where the mass of inner planet is ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$). The numerical results shown in Figure 8 can directly be compared with the schematic presented in Figure 6.
We find that the simulations agree pretty well with the analytical estimates. No planets lead resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (upon the green line) and slow migration (left of the brown line) criteria break. When those trapping criteria are satisfied and as $\tau_{\rm a}/\tau_{\rm e}$ decreases, the simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes. These results are indeed separated by the two analytical criteria proposed in Sect. 2 (yellow and purple lines).
However, there is one point in Figure 8 that simulations do not fully match the analysis, occurring at $\tau_{a}{\sim}10^{5}\ {\rm yr}$ and $\tau_{a}/\tau_{e}{\sim}3\times 10^{2}$. The theory predicts the planets in overstable trap regime whereas in simulations planets actually leave the resonance. We speculate that the reason is that as migration becomes faster, the perturbations intensify and become nonlinear, resulting in further increasing oscillation amplitudes around the equilibrium value. Consequently, the system becomes more likely to cross the separatrix and escape resonance.
Figure 8 demonstrates simulation results from the case when the mass of inner planet is ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), serving as a comparison to the schematic illustrated in Figure 6. Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied. When these criteria are fulfilled, all traps correspond to stable trap (green dots).
We note that the no-trap regime occupies a large parameter space than theoretically expected. This is evident at $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$, where the trapping criteria are fulfilled but the planets are not in resonance. This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq. 6). When the eccentricity damping is considered, the interaction between the planets is further suppressed (lower $e$ in left-hand side of Eq. 3), leading to a longer migration timescale (right-hand side of Eq. 3). Consequently, in reality planets can bypass the resonance with slower migration rates compared to the simplified estimation from Eq. 6.
In conclusion, our analytical predictions align well with simulations in the circumstances of planet pairs with extreme mass ratios.
3.3 Comparison at arbitrary mass ratios
Notably, Huang & Ormel (2023) derived the weak damping trapping criterion at the small mass ratio limit of $q{\ll}1$. Goldreich & Schlichting (2014) investigated when the trap becomes overstability and/or escape in the same limit. In this work we derived the general criteria for trap and stability that are well suited to the arbitrary mass ratio (Eqs. 13 and 24). In this subsection we show the simulation comparison, whereas the detailed derivation of the $q$-dependent overstabity and escape criteria is provided in Appendix B.
Figure 9 illustrates the simulation results for the equal-mass planet case ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$) and compares our criteria with those from Huang & Ormel (2023) and Goldreich & Schlichting (2014). The solid lines represent our analytical criteria, dashed lines correspond to criteria from previous studies, and colored dots depict the simulation results. We find that the new criteria clearly depict a better agreement with numerical results.
Figure 10: The plots illustrate the dependence of resonance trapping criteria on the semi-major axis (panel a) and the stellar mass (panel b) for a two-planet system, based on the minimum solar nebular model (MMSN). The planet masses are $m_{\rm i}=1\ M_{\oplus}$ and $m_{\rm o}=10\ M_{\oplus}$. In panel a, the stellar mass is fixed at $M_{\star}=1\ M_{\odot}$, while in panel b, the semi-major axis of the inner planet is fixed at $a_{\rm i}=1\ au$. The brown line represents the slow migration trapping criterion (Eq. 8), below which migration is too fast for resonance capture. Regions above the brown line indicate where capture is possible. The color gradient represents the strongest eccentricity damping $\tau_{a0}/\tau_{e0}$ that allows resonant capture (Eq. 11), with lighter colors meaning the planets can be in resonances with stronger eccentricity damping. In panel (a), as the semi-major axis increases, the critical value of $\tau_{a0}$ remains constant, while $\tau_{a0}/\tau_{e0}$ slightly increases. This suggests that planets are more easily trapped in resonances under stronger eccentricity damping as the semi-major axis increases. In panel (b), as the stellar mass decreases, the critical value of $\tau_{a0}$ decreases, and $\tau_{a0}/\tau_{e0}$ increases. This indicates that planets are more easily trapped in resonances under stronger eccentricity damping and faster migration as the stellar mass decreases.
4 Discussion
In this section, we connect our criteria to disk parameters in Sec. 4.1 and analyze how the criteria depend on the planet location and the mass of stellar host. Additionally, we investigate resonance trapping in the context of divergent migration in Sec. 4.2 and discuss several caveats of our work in Sec. 4.3.
4.1 Dependence on semi-major axis and stellar mass
Here we explore how the resonant trapping conditions depend on the radial distance and stellar mass. The detailed type I migration formulas for low-mass planets are given by $\tau_{a}^{-1}{\sim}\mu\mu_{\rm d}h^{-2}n$ and $\tau_{e}^{-1}\sim\mu\mu_{\rm d}h^{-4}n$, where $\mu_{\rm d}{=}\Sigma_{\rm d}a^{2}/M_{\rm*}$ is the mass ratio between disk and star, and $\Sigma_{\rm d}$ and $h$ are the disk surface density and aspect ratio, respectively.
Following the minimum mass solar nebula (MMSN, Hayashi 1981) model, the disk surface density and aspect ratio can be written as
$\Sigma_{\rm d}=$ $1700\left(\frac{a}{\rm 1\ au}\right)^{-3/2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)\rm\ g/cm^{2},$ (28)
$h=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{L_{\star}}{1\ L_{\odot}}\right)^{1/8}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/2}$
$=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/4},$ (29)
the latter of which $L_{\star}{\propto}M_{\star}^{2}$ is assumed for the pre-main sequence stars. By substituting $\Sigma_{\rm d}$ and $h$ into $\tau_{a}$ and $\tau_{e}$, we obtain their dependence on the semi-major axis and stellar mass as
$\tau_{a}={\tau_{a0}}\left(\frac{a}{\rm 1\ au}\right)^{3/2},\ \ \tau_{e}={\tau_{e0}}\left(\frac{a}{\rm 1\ au}\right)^{2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{1/2},$ (30)
where $\tau_{a0}$ and $\tau_{e0}$ represent the values of $\tau_{a}$ and $\tau_{e}$ at $1$ au around a solar-mass host. By inserting $\tau_{a0}$ and $\tau_{e0}$ into Eq. (13), we can separate the dependence of resonance trapping on semi-major axis and stellar mass.
Figure 10 shows the a-dependency for resonance trapping, where $M_{\star}{=}1\ M_{\odot}$ and $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$. The region below the brown line (slow migration criterion) indicates that migration is too fast for resonance capture, while the region above suggests that capture is possible, which depends on $\tau_{a0}/\tau_{e0}$. The color in Figure 10 refers to the critical $\left(\tau_{a0}/\tau_{e0}\right)$ that allows resonant capture. A lighter color means the planets can be in resonances with higher $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ (stronger e-damping). As the semi-major axis increases, $(\tau_{a0})_{\rm slow,mig}$ remains constant and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ slightly increases. This means that planets would be more easily trapped at resonances under stronger e-damping as $a$ increases. Nevertheless, the capture reveals an overall weak semi-major axis dependence.
Figure 10 shows $M_{\star}$-dependency for resonance trapping. We fix the inner planet at $1\ au$ and vary with the stellar mass. From the figure, as the stellar mass decreases, $(\tau_{a0})_{\rm slow,mig}$ decreases, and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ increases. This implies that planets can be trapped under a wider range of $\tau_{a}$ and $\tau_{e}$ parameter space. In other words, planets are more likely to be capture into resonance as the stellar mass decreases.
Based on the MMSN model, we find that planets are more likely to get capture into resonance with increasing semi-major axis and/or decreasing stellar mass. However, we note that these results are considerably sensitive to the adopted disk model. The interpretation should be taken with caution.
4.2 Divergent migration
@@
We have shown in the previous sections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping. Literature theoretical work suggested that planets cannot be captured in resonance when the direction of their migration is divergent (Wong & Lee, 2024). Based on our analysis, here we also provide a simple argument that why resonance trapping cannot occur through divergent migration.
@@
where the relative orbital decay term ($-\tau_{a}^{-1}$) is negative compared to Eq. (3). In order to derive equilibrium eccentricity, we repeat the procedure outlined in Sec. 2.1 by substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$. Then Eq. (31) reads
$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{2\tau_{a}}.$ (32)
@@
Noticeably, the left-hand side of this equation is positive, whereas the right-hand side is negative. This indicates that there is no equilibrium solution in the divergent migration circumstance. We also conducted numerical simulations and consistently found that even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
4.3 Caveats

### v2

@@ 2.3 A unified picture
For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm m}/\tau_{e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime). As the relative damping weakens (Eq. 22), the fixed point becomes unstable and the trajectory of the system begins to oscillate around the fixed point. It ultimately settles into a stable limit cycle (overstable trap). The limit cycle expands as $\tau_{\rm m}/\tau_{e}$ continues to decrease. When the trajectory crosses the separatrix (Eq. 23), the system becomes unstable. In this regime, the oscillations grow over time and planets eventually leave resonance (escape). In contrast, for the case of a small outer planet shown in Figure 6, all traps are stable.
Figure 7: Simulations of the 22:11 MMR capture and stability for two-planet systems with an inner to outer planet mass ratio $q$ of $0.1$. Blue: planets bypass resonance due to rapid migration or strong eccentricity damping. Green: stable resonance capture. Yellow: overstable resonant capture. Purple: Resonance escape after temporary capture. The simulations are consistent with the analysis outlined in Figure 6.
Figure 8: Same as Figure 7 but with an inner to outer planet mass ratio $q$ of $10$.
Figure 9: Same as Figure 7 but with equal-mass of q=1q{=}1. The solid lines are the criteria derived in this study, whereas the dashed lines represent the trapping criteria of Batygin (2015) (their Eq. 43), and the stability and escape criteria of Deck & Batygin (2015) (their Eqs. 15 and 21).
2.4 Resonance crossing through divergent migration
@@
We have shown in the previous subsections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping. Literature theoretical work suggested that planets directly cross the resonance when the direction of their migration is divergent due to separatrix crossing (Henrard & Lemaitre, 1983; Murray & Dermott, 1999; Wong & Lee, 2024). Here we also explain that divergent migration cannot lead to resonance capture in our framework.
@@
where the relative orbital decay term ($\tau_{\rm m}^{-1}$) is negative compared to Eq. (3). In order to derive equilibrium eccentricity, we repeat the procedure outlined in Sec. 2.1 by substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$. Then Eq. (27) reads
$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{\tau_{\rm m}}.$ (28)
@@
Noticeably, as long as $\tau_{\rm m}$ and $\tau_{\rm e}$ have the same sign, the left-hand side of this equation has the opposite sign as the right-hand side, and therefore the equality never holds. This indicates that there is no equilibrium solution in the divergent migration circumstance, and even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
3 Simulations
In this section, we conduct numerical simulations to verify the aforementioned analytical criteria. The numerical setup is described in Section 3.1. We first test the 22:11 resonance for planets with extreme mass ratios in Section 3.2, and then extend the analysis for planets with arbitrary mass ratios in Section 3.3. Other first-order resonances are investigated in Section 3.4.
3.1 Numerical setup
We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012). In simulations we also consider planets in coplanar orbits. The accelerations for angular momentum and eccentricity damping are given by (Papaloizou & Larwood, 2000; Cresswell & Nelson, 2006, 2008; Pichierri et al., 2024)
${a_{\rm m}}=-\frac{\textbf{v}_{\rm pl}}{\tau_{\rm m}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}_{\rm pl}\cdot\textbf{r}_{\rm pl})\cdot\textbf{r}_{\rm pl}}{r_{\rm pl}^{2}\tau_{e}},$ (29)
which are implemented through REBOUNDx (Tamayo et al., 2020).
Initially, the inner planet is placed at $1$ au and the orbit of the outer planet is slightly away from the desired first-order resonance ( e.g., at $1.7$ au for the 22:11 resonance). Both planets have circular orbits with random phase angles. To mimic convergent migration, only the outer planet undergoes inward migration on a timescale of $\tau_{\rm m}$, and the eccentricities of both planets are damped on a timescale of $\tau_{e}$. Simulations are terminated at $t{=}2\tau_{\rm m}$, and the outcomes can be classified in different regimes, as described in Section 2.3.
We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, q=0.1q{=}0.1), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, q=10q{=}10), and equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, q=1q{=}1). We vary $\tau_{\rm m}$ and $\tau_{e}$ in each simulation.
3.2 Systems at two extreme mass ratios
Figure 8 shows the simulations for the planet pair with the inner planet mass ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, q=0.1q{=}0.1), the results of which can be compared with the schematic presented in Figure 6.
We find that the simulations agree pretty well with the analytical estimates. No planets lead the resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (green line) and slow migration (brown line) criteria break. When those trapping criteria are satisfied and as $\tau_{\rm m}/\tau_{e}$ decreases, the simulations depict a transition from the stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes. These results are indeed separated by the two analytical criteria proposed in Sect. 2 (yellow and purple lines).
The simulations do not fully match the analysis at $\tau_{\rm m}{\sim}2\ \times 10^{5}\ {\rm yr}$ and $\tau_{\rm m}/\tau_{e}{\sim}6\times 10^{2}$ in Figure 8. The theory predicts the planets in an overstable trap regime whereas in simulations planets actually leave the resonance. We hypothesize that the increased relative migration rate leads to more intense and nonlinear perturbations. This results in greater oscillation amplitudes around the equilibrium point. As such, the system has a higher probability of crossing the separatrix and escaping resonance.
Figure 8 demonstrates simulations for the planet pair with the inner planet mass ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, q=10q{=}10), offering a comparison to the schematic illustrated in Figure 6. Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied. When these criteria are fulfilled, all traps correspond to stable traps (green dots).
We note that the no-trap regime occupies a slightly larger parameter space than theoretically expected. For instance, at $\tau_{\rm m}{\sim}4{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm m}/\tau_{e}{\sim}4\times 10^{2}$, while the trapping criteria are fulfilled, the planets are not in resonance. This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq. 6). When the eccentricity damping is incorporated, the interaction between the planets is further suppressed—owing to reduced eccentricities in the left-hand side of Eq. 3—thereby prolonging the migration timescale on the right-hand side of Eq. 3. Consequently, planets in reality bypass the resonance with slower migration rates compared to the simplified estimation from Eq. 6.
Overall, we find that the analytical criteria align well with simulations in the circumstances of the planet pairs with extreme mass ratios.
3.3 systems at arbitrary mass ratios
The resonance dynamics has been extensively investigated in literature studies. Batygin (2015) proposed an adiabatic capture criterion for systems without eccentricity damping, linking the resonant libration time to the migration time. Subsequently, Huang & Ormel (2023) derived a weak eccentricity damping criterion for planet pairs with extreme mass ratios ($q{\ll}1$), while Batygin & Petit (2023) formulated a generalized criterion applicable to arbitrary mass ratios. Regarding the post-capture evolution, Goldreich & Schlichting (2014) first proposed the concept of overstability and derived the corresponding analytic criteria for restricted three-body systems. Deck & Batygin (2015) further extended the analysis for planet pairs with arbitrary mass ratios.
In this subsection, we conduct numerical simulations to validate our theoretical criteria and compare our results with those from literature studies. The detailed derivation of the $q$-dependent overstability and escape criteria are provided in Appendix B. As a point of comparison, our weak damping criterion is equivalent to that obtained by Batygin & Petit (2023), while our other criteria exhibit the same $\mu$-scaling as those in previous works but differ in their numerical prefactors: $(\tau_{\rm m})_{\rm slow,mig}{\sim}\mu^{-4/3}$, $(\tau_{\rm m}/\tau_{e})_{\rm overstab},(\tau_{\rm m}/\tau_{e})_{\rm esc}{\sim}\mu^{-2/3}$. Therefore, we compare our slow migration criterion with the adiabatic criterion in Batygin (2015) and our stability and escape criteria with those in Deck & Batygin (2015). Additionally, we perform simulations to investigate the $q$-dependence of our overstability and escape criteria, as detailed in Appendix B.
Figure 9 illustrates the simulations for the planet pair of equal-mass ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, q=1q{=}1). The colored symbol depicts the simulation results, the solid lines represent our analytical criteria (Eq. 24), and the dashed lines correspond to the criteria from Batygin (2015) (their Eq 43) and Deck & Batygin (2015) (their Eqs 15 and 21). As can be seen, both the criteria derived in this work and those from previous studies match the numerical results relatively well. A more detailed comparison that accounts for the dependence on the mass ratio $q$ is presented in Appendix B.
3.4 Systems at arbitrary j:j-1 MMR
In previous sections we focus on the circumstances of the 22:11 resonance as a representative case. Since our analytical framework can be applied to other first-order $j$:j1j-1 MMRs, we also perform such numerical verification as follows.
Figure 10 shows the resonant states in which the systems finally settle for equal-mass planet pairs ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$). The colored dots depict the numerical results, the green lines represent the trapping criteria, while the purple line represents the escape criterion. Note that for equal-mass planet pairs, our criteria predict that all traps are stable (no overstable trap and escape regimes) at $j{\geq}3$ (Eq. 24). Therefore, we only exhibit the escape criterion for the 22:11 MMR in Figure 10.
Our criteria show good agreement with simulations. As the disk effects become stronger (e.g., $\tau_{\rm m}$ decreases and $\tau_{\rm m}/\tau_{e}$ increases), planets can bypass the lower $j$:j1j-1 MMRs and evolve into more compact orbits.
Figure 10: Simulations of $j$:j1j-1 MMR capture and stability for two-planet systems with equal masses. Color dots represent systems’ final resonant states. Green lines show $j$:j1j-1 trapping criteria, while purple indicates the 22:11 escape criterion.
4 Discussion